OpenStudy (yuki):

Can someone help me out with linear algebra? I need to find the matrix of transformation for the following.

OpenStudy (yuki):

\[T(\left[\begin{matrix}a & b \\ c & d\end{matrix}\right]) = \left[\begin{matrix}10a-9b & 4a-9b \\ -2c+5d & c+2d\end{matrix}\right]\]

OpenStudy (yuki):

you guys are just bunch of idiots who can't even solve one college math problem